OpenStudy (anonymous):

Find an equation of the line that passes through the given point and has the given slope. Express your answer in slope– intercept form. (8,0), m = -3

OpenStudy (anonymous):

sorry that was slope point form, the slope intercept form is y=mx+b, so y=-3x+24 I just put it in slope point form then change it to y intercept y-y1=m(x-x1)

OpenStudy (tkhunny):

Have you considered the Point-Slope form to get started?

OpenStudy (anonymous):

y = mx + b (solving for b, the y intercept) slope(m) = -3 (8,0) x = 8 and y = 0 now we sub y = mx + b 0 = -3(8) + b 0 = -24 + b 24 = b your line is : y = -3x + 24 I agree with your first post :)

OpenStudy (anonymous):

So many ways to solve a problem :)